The perimeter of this block of land is 72 m. Find:
a the length of the block
b the area of the block.

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

you don't have enough information to determine the length of the block or the area of the block.

the area of the block will differ depending on the length of the block and the width of the block.

the length of the block requires you to know the width of the block before determining what it is.

here's some examples:

72 is the perimeter.
you are given that.

the perimeter is determined by the equation 2L + 2W = perimeter.

you have 2L + 2W = 72

if you let W = 8, then 2W = 16 and the formula becomes:
2L + 16 = 72
subtract 16 from both sides of this equation to get:
2L = 56
divide this equation by 2 to get:
L = 28

your dimensions are:
L = 28
W = 8

the area is L * W which equals 28 * 8 which equals 224 square meters.

now, assume W = 16.
this means that 2W = 32.
this means that 2L + 2W = 72 becomes 2L + 32 = 72
subtract 32 from both sides of that equation to get:
2L = 40
divide both sides of that equation to get
L = 20

your dimensions are:

L = 20
W = 16

the area is now L * W = 20 * 16 = 320 square units.

Given that the perimeter is 72 meters, the length is dependent on the width and the area is dependent on the length times the width.

without knowing the width, you can have many possible lengths depending on the dimension of the width that you choose.

here's a table that shows you the possible lengths and widths with a perimeter of 72.

these are all integers in multiples of 4.
if you assume real numbers that could include fractional parts of a whole number, then the list can be endless.

Perimeter
	Length	Width	Area
72	4	68	272
72	8	64	512
72	12	60	720
72	16	56	896
72	20	52	1040
72	24	48	1152
72	28	44	1232
72	32	40	1280
72	36	36	1296
72	40	32	1280
72	44	28	1232
72	48	24	1152
72	52	20	1040
72	56	16	896
72	60	12	720
72	64	8	512
72	68	4	272



Answer by ikleyn(53427)

.
The perimeter of this block of land is 72 m. Find:
(a) the length of the block
(b) the area of the block.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


This problem is incomplete and can not be solved/answered.

Before submitting to this forum, read your posts attentively
and prepare them as accurately as it should be done.